Find the area under the normal curve to the left of z= -0.28 plus the area under the normal curve to the right of z =1.40.

The combined area is?
Answer #1

Using standard normal z-table we get,

P(Z < -0.28) = 1 - P(Z < 0.28) = 1 - 0.6103 = 0.3897

and

P(Z > 1.40) = 1 - P(Z < 1.40) = 1 - 0.9192 = 0.08078

Therefore, we get,

P(Z < -0.28) + P(Z > 1.40) = 0.3897 + 0.0808 = 0.4705

=> The combined area is 0.4705